4) Machine A costs $35,000, lasts 3 years and has a salvage value of $17,500. Machine B costs
$25,000, lasts 3 years and has a salvage value of $6,500. What is the ROR on the incremental
investment of Machine A over Machine B? What range of MARR would make you prefer
Machine B?
Answer #1

ANSWER:

a) In order to find the incremental ror we will equate the present worth to zero.

pw(a - b) = cost of machine a - cost of machine b + salvage value of machine a - salvage value of machine b(p/f,i,n)

0 = -35,000 - (-25,000) + 17,500 - 6,500 (p/f,i,3)

0 = -10,000 + 11,000(p/f,i,3)

10,000 = 11,000(p/f,i,3)

10,000 / 11,000 = (p/f,i,3)

0.909 = (p/f,i,3)

solving via trial and error we get that ror is 3.23%

b) MARR over 3.23% and beyond will make me prefer machine b as in that case the incremental npv will be more.